Answer the following with full solution.


1. ) A shot is project 40° from the horizontal at the rate of 30 ft/s. (a) Find the maximum height (b) the maximum range of projectile.

2. ) A drop hammer of 1 ton weight capacity is projected downward by a 12 inch diameter cylinder. At 100 psi air pressure, what is the impact velocity if the stroke is 28 inches.

3.) A projectile is fired up the inclined plane at an initial velocity of 15 m/s. The plane is making an angle of 30° from the horizontal. If the projectile was fired at 30° from the incline, compute the maximum height z measured perpendicular to the incline that is reached by the projectile. Neglect air resistance.


1
Expert's answer
2022-03-27T18:20:46-0400

1.

"h=\frac{v^2\sin^2\alpha}{2g}=5.78~ft,"

"l=\frac{v^2\sin 2\alpha}g=27.57~ft,"

2.

"\frac{mv^2}2=(mg+pS)h,"

"v=\sqrt{2h(g+\frac{pS}m)}=31.6~\frac{ft}s,"

3.

"h=\frac{v^2\sin^2\alpha}{2g}\cos\beta=2.5~m."